Compute the value of the ending inventory. Costs increase 5% every quarter over the beginning cost. 6000 units are sold durin it is a Management system control question
plz show how to get the answer
Answer #1
Qty Cost
B 1500 200000
Q1 1500 210000
Q2 1500 220000
Q3 1500 230000
Q4 1500 240000
TOTAL 7500 1100000

6000 units sold as given in question. So balance ie 7500-6000 = 1500 in stock at end.

a) FIFO: Oldest inventory in hand will be first sold, then the succeeding inventory will be sold

COGS 6000 200000+210000+220000+230000 = 860000
Ending 1500 240000

b) LIFO: Newest inventory in hand will be first sold, then the preceding inventory will follow:

COGS 6000 240000+230000+220000= 900000
Ending 1500 200000

Please comment for any further help on this solution.